OpenStudy (anonymous):

Solve : -2 + -2(x - 3) = 8

OpenStudy (lgbasallote):

first...add 2 to both sides..what do you get?

OpenStudy (anonymous):

10

OpenStudy (anonymous):

expand the brackets -2 -2x +6 = 8 Simplify add -2 + 6 =4 4 -2x =8 put numbers on one side and letters on the other. so minus 4 both sides -2x =4 You can multiply by minus 1 to make the x term positive 2x = -4 Divide by 2 x = -2

OpenStudy (anonymous):

let me know if that helped

OpenStudy (anonymous):

these are the answer choises.x=3x=0x=4/3x=5